OpenStudy (anonymous):

hi. how do you find n mathematically for a question like: find the sum of the finite geometric series -5-10-20-40-...-2560

OpenStudy (anonymous):

There are formulas for all types of geometric progression Easily available online

OpenStudy (anonymous):

well the formula i have is s=a1(1-r^n)/1-r, but i don't understand how to find n...

OpenStudy (amistre64):

if you know r and s .... its just algebra

OpenStudy (amistre64):

even if you dont know them ... its still algebra :)

OpenStudy (anonymous):

you don't know s. you're trying to find the sum

OpenStudy (amistre64):

\[S_n=a_1\frac{1-r^n}{1-r}\] \[S_n\frac{1-r}{a_1}=1-r^n\] \[r^n=1-S_n\frac{1-r}{a_1}\] \[log~r^n=log\left(1-S_n\frac{1-r}{a_1}\right)\] \[n~log~r=log\left(1-S_n\frac{1-r}{a_1}\right)\] \[n=\frac{log\left(1-S_n\frac{1-r}{a_1}\right)}{log(r)}\]

OpenStudy (amistre64):

so your trying to find out the number of terms in order to find the sum, thats a different thing

OpenStudy (anonymous):

yeah im trying to find the sum

OpenStudy (amistre64):

you have to use the explicit formula, not the sumation formula \[a_n=a_1~r^{n-1}\] \[2560=5~(2)^{n-1}\]

OpenStudy (anonymous):

so from there you solve for n?

OpenStudy (amistre64):

yep

OpenStudy (amistre64):

looks like n=2560(2/5) to me

OpenStudy (anonymous):

ok ill do it that way thanks a ton for your help

OpenStudy (amistre64):

youre welcome.